OpenStudy (anonymous):

What is the simplest form of the radical expression: 3^4(sqrt2x) minus 6^4(sqrt2x)? Please show me how you did it too. I don't get it at all please help!!!!

OpenStudy (anonymous):

\[3^{4}\sqrt{2x}-6^{4}\sqrt{2x}\]

OpenStudy (anonymous):

I will medal+fan if you answer!! Please help me!

OpenStudy (anonymous):

\[\begin{align*}3^4\sqrt{2x}-6^4\sqrt{2x}&=3^4\sqrt{2x}-(2\times3)^4\sqrt{2x}\\\\ &=3^4\sqrt{2x}-2^4\times3^4\sqrt{2x}\\\\ &=3^4\sqrt{2x}\left(1-2^4\right)\\\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

Sooo what would it equal?